# Revealing Multimodal Contrastive Representation Learning through Latent Partial Causal Models

### v1

@@ 5 Experiments
Table 1: Assessing identifiability up to linear transformations under varying assumptions. The first row corresponds to a setting that matches our assumptions in theorem 3.1, while the others show results for violated assumptions. S: Space, Sp: Sphere, U: Uniform, v: vMF ($k=1$), L: Laplace ($\lambda=0.05$), N: Normal ($\delta=0.05$), B: Box, Un: Unbounded, G: GenNorm($\beta=3$)
@@
Table 2: Assessing identifiability up to permutation under varying assumptions. The first two rows correspond to a setting that matches our assumptions in theorem 3.2, while the others show results for violated assumptions. S: Space, Sp: Sphere, U: Uniform, v: vMF ($k=1$), L: Laplace ($\lambda=0.05$), N: Normal ($\delta=0.05$), B: Box, Un: Unbounded, G: GenNorm ($\beta=3,\lambda=0.05$)
@@
To test for identifiability up to linear transformations formalized by Theorem 3.1, we fit a linear regression model between the ground-truth $\mathbf{z}_{x}$ and recovered $\mathbf{\hat{z}}_{x}$ and report the coefficient of determination ($R^{2}$). Further, to test for identifiability up to permutations formalized by theorem 3.2, we employ the mean correlation coefficient (MCC). The first row in Table 1 and the first two rows in Table 2, corresponding to the setting where the assumptions are satisfied, verify the identifiability results on hypersphere and convex bodies, respectively. Our empirical investigations have yielded a critical insight: discrepancies in the assumptions concerning marginal and conditional distributions, as well as the nature of the spaces (hypersphere and convex body), do not significantly impact performance. This robustness is demonstrated by the results detailed in Table 1 for the hypersphere space and Table 2 for convex bodies. This observation is similar to reports from studies conducted in the context of single-model context (Zimmermann et al., 2021). This observation might be attributed to the fact that the loss function described in Eq. 2 predominantly relies on the computation of expectations, inherently allowing for a wide range of approximations. If we can approximate the expectation calculations consistently across various distributions and spaces, it is reasonable to expect that the identifiability results remain well within acceptable bounds.
@@
Informed by our identifiability results and the empirical evidence presented earlier, we have grounds to claim that the pre-trained CLIP model possesses disentanglement ability. To substantiate this, as discussed in Section 4, we first extract features from the pre-trained CLIP model and then apply FastICA to these features to achieve final representations. We expect these final representations to exhibit clear signs of disentanglement. To validate this, we proceed to train a decoder that reconstructs observational data using these extracted representations. We implement the above process on the CelebA face dataset (Liu et al., 2015), which has been explored by previous studies to learn disentangled representations (Kim & Mnih, 2018; Chen et al., 2018).
Figure 2 illustrates the effectiveness of our method through latent space traversals. Specifically, it visualizes changes in reconstructions as we traverse one dimension of the latent space at a time, showcasing 4 out of 16 attributes uncovered by our approach. Our method yields competitive results when compared with specialized techniques for learning disentangled representations, such as FactorVAE(Kim & Mnih, 2018) and $\beta$-TCVAE (Chen et al., 2018). As reported, FactorVAE identified 8 disentangled attributes and $\beta$-TCVAE reported 15, our method successfully discerns 16 distinct disentangled representations. Additional results are available in Appendix A.6. This achievement not only underscores the effectiveness of our method and validates our identifiability results, but also offers new perspectives and insights into learning disentangled representations by pre-trained CLIP.
(a) Smile
(b) Gender w/ Mustache
(c) Glasses
(d) Face Size
Figure 2: Disentangled Representations learned by combining pre-train CLIP and FastICA. The proposed method obtains 16 disentangled representations, refer to Appendix A.6 for more results.
Table 3: Quantitative results for 2-shot transfer learning and domain generalization by different methods. Lin. P. (Linear Probe).
@@
RN50 Lin. P. 31.95 26.48 8.41 20.74 7.44 15.77
Lin. P. w/ FastICA 34.06 28.74 8.37 21.72 10.15 17.25
Lin. P. w/ PCA and FastICA 34.12 28.68 11.55 25.57 10.15 18.99
RN101 Lin. P. 37.64 31.45 13.71 31.09 11.85 20.03
Lin. P. w/ FastICA 39.58 33.15 13.49 30.29 14.77 22.93
Lin. P. w/ PCA and FastICA 39.86 33.58 17.93 35.48 14.20 25.29
ViT32 Lin. P. 38.23 32.00 16.17 33.67 12.88 23.68
Lin. P. w/ FastICA 40.21 33.97 16.54 34.79 15.72 25.26
Lin. P. w/ PCA and FastICA 39.34 33.44 19.02 36.98 14.69 26.03
ViT16 Lin. P. 44.97 38.11 22.06 43.86 25.99 32.51
Lin. P. w/ FastICA 45.52 39.38 22.55 45.33 30.47 34.43
Lin. P. w/ PCA and FastICA 46.57 40.66 26.67 49.69 31.48 37.13
Source Target (ImageNet-)
Encoders Methods ImageNet V2 Sketch R A Avg.
RN50 Lin. P. 41.34 33.67 11.55 26.27 9.67 20.29
Lin. P. w/ FastICA 44.10 36.07 12.75 30.15 11.64 22.65
Lin. P. w/ PCA and FastICA 42.86 35.38 12.29 28.81 9.79 21.57
RN101 Lin. P. 48.23 39.53 18.80 38.10 14.32 27.69
Lin. P. w/ FastICA 49.43 41.02 17.49 39.33 15.25 28.27
Lin. P. w/ PCA and FastICA 49.01 40.25 19.26 39.71 14.75 28.49
ViT32 Lin. P. 47.82 39.53 21.51 40.94 15.99 29.49
Lin. P. w/ FastICA 49.43 40.66 22.66 41.78 16.41 30.38
Lin. P. w/ PCA and FastICA 49.48 41.09 23.72 43.48 16.77 31.27
ViT16 Lin. P. 54.30 46.06 27.58 50.76 29.24 38.41
Lin. P. w/ FastICA 56.65 48.18 28.27 55.50 33.39 41.33
Lin. P. w/ PCA and FastICA 56.16 47.46 30.21 55.49 31.71 41.22
Few-shot learning and domain generalization on ImageNet-type data
Broadly, the objective of disentangled representations is to learn features that lend themselves to be easily and robustly transferred to downstream tasks. This implies that disentangled representations should inherently possess a satisfactory capability for few-shot learning and demonstrate robustness against distribution shifts. Therefore, we focus on tasks involving few-shot learning and domain generalization, to validate our identifiability results and the efficacy of the proposed methods. To achieve this, we first obtain representations of a limited set of labeled samples. This is done either by utilizing the pre-trained CLIP model followed by FastICA (labeled as Linear Probe with FastICA, applied within the hypersphere space) or by employing the pre-trained CLIP model in conjunction with PCA and FastICA (labeled as Linear Probe with PCA and FastICA, aligning with convex bodies). These extracted representations, along with their labels, are used to train a linear classifier. We train the proposed methods on ImageNet (Deng et al., 2009) with limited samples to evaluate their performance for few-shot learning, and also conduct evaluations on ImageNet-V2 (Recht et al., 2019), ImageNet-Sketch (Wang et al., 2019), ImageNet-R (Hendrycks et al., 2021a), and ImageNet-A (Hendrycks et al., 2021b) for demonstrating the robustness to distribution shift.
Tables 3 and 4 present the performance metrics of the proposed methods in few-shot learning scenarios (as shown in the ‘SOURCE’ column) and in the context of distribution shift (as indicated in the ‘TARGET’ columns). An analysis of the data for the ‘SOURCE’ column in these tables reveals that the proposed methods outperform the baseline approach of training a linear classifier with features directly obtained from pre-trained CLIP, known as Linear Probe. This superior performance underscores the enhanced adaptability of our proposed methods, particularly in tasks requiring rapid learning from limited data. Furthermore, observations from the ‘TARGET’ column demonstrate that our proposed methods also surpass the Linear Probe approach in terms of distribution shift, which affirms the robustness of our methods. See Appendix A.7 for more results.
@@ 6 Conclusion
In this study, we demonstrate that multimodal contrastive representation learning can effectively identify latent coupled variables up to a linear transformation on the hypersphere and up to permutation on convex bodies. This insight significantly contributes to our understanding of the success factors behind multimodal contrastive representation learning. Building on these theoretical findings, we investigate the disentanglement ability of pre-trained multimodal models such as CLIP, each tailored to different spaces. Our experiments with synthetic data not only corroborate our theoretical assertions but also empirically show that the identifiability results remain valid even when various assumptions are not met. Further, our experiments on the CelebA dataset, few-shot learning, and domain generalization, highlight the potential for learning disentangled representations by pre-trained CLIP. While our findings are promising, the pursuit of even more effective methods to unlock and harness the full potential of disentanglement in pre-trained CLIP models remains a valuable and ongoing endeavor.
7 Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Appendix
A.1 Related Work
@@
Multi-modal contrastive representation learning, driven by underlying co-occurrence patterns across various modalities, aims to coalesce inputs from these diverse sources into a cohesive representation space. This is typically achieved using a symmetric version of the standard contrastive loss (Oord et al., 2018; Gutmann & Hyvärinen, 2010), a method designed to align accurate pairings while distinguishing incorrect ones (He & Peng, 2017; Radford et al., 2021). Although this approach has proven successful in a range of downstream tasks (Radford et al., 2021; Zhou et al., 2022a, b; Lüddecke & Ecker, 2022; Ban & Dong, 2022), there remains a gap in our comprehensive theoretical and empirical understanding of the representations it learns. Recently, there has been a growing interest in exploring multi-modal contrastive learning from various perspectives. For instance, the study by Liang et al. (2022) provides insights into the modality gap inherent in multi-modal contrastive learning. Similarly, the research presented by Nakada et al. (2023) establishes a link between general multimodal contrastive loss and SVD analysis. Additionally, Huang et al. (2021) posits that learning with multiple modalities can lead to a reduced population risk compared to using a subset of these modalities. Diverging from these approaches, our work delves into multi-modal contrastive representation learning by examining its connection with generative models.
@@
Past research has sought to comprehend the representations derived from standard single-modality contrastive learning, examining them through the lens of alignment and uniformity (Wang & Isola, 2020), showing guarantees on the performance of the learned representations on the average classification task (Saunshi et al., 2019), or in terms of the identifiability of latent variables (Zimmermann et al., 2021; Von Kügelgen et al., 2021). Building on these foundations, our work takes a foreword step. We demonstrate that multi-modal contrastive learning can identify latent coupled variables, extending the insights from previous studies into the realm of multi-modality.
@@
Very recently, several studies have emerged, focusing on the intersection between the identifiability of latent variables and modal contrastive representation learning (Daunhawer et al., 2023; Gresele et al., 2020; Yao et al., 2023). Some of these works, however, have only achieved partial identifiability of coupled variables (Daunhawer et al., 2023; Yao et al., 2023), specifically identifying latent content variables but not latent style variables. In contrast, our work achieves comprehensive identifiability results for all latent coupled variables, offering a deeper level of understanding. Our research also diverges from the approach taken in (Gresele et al., 2020) in two key ways: Firstly, we model co-occurrence patterns using conditional distributions, whereas the latter utilizes identical variables for this purpose. Secondly, while (Gresele et al., 2020) relies on the premise that the mapping from the latent space to observations must be constrained by component-wise corrupters to ensure identifiability, our findings do not necessitate such constraints.
@@
Nonlinear Independent Component Analysis (ICA) aims to unravel latent independent variables from observational data that has been subject to a nonlinear mixture of these latent factors. However, as pointed out in the seminal work by Hyvärinen et al. (Hyvärinen & Pajunen, 1999), solving this problem is generally infeasible without specific underlying assumptions. A prominent direction in contemporary research leverages the concept of distributional changes in latent variables, which leads to the creation of multi-domain observational data. This approach has been extensively explored and developed in a series of studies (Hyvarinen & Morioka, 2016, 2017; Hyvarinen et al., 2019; Khemakhem et al., 2020), each contributing to a deeper understanding and more refined methodologies in the field of Nonlinear ICA. We build upon this body of research by incorporating co-occurrence patterns observed across multiple modalities. It is important to note the distinct difference between multi-domain and multi-modal approaches. The former typically implies a consistent mapping from the latent space to the observational space across all domains, whereas the latter accommodates different mappings for each modality. Additionally, while multi-domain approaches generally assume a totally shared latent variables across all domains, multi-modal methods allow for the existence of modality-specific latent variables.
A.2 The Proof of Theorem 2.1
@@
(The asymptotics of $\mathcal{L}$) For fixed $\tau>0$, as the sample size $N\rightarrow\infty$, the (normalized) symmetric contrastive loss converges to
$\lim_{N\rightarrow\infty}\mathcal{L}-2\log N=$ $-2\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{x},\mathbf{t})\sim p(\mathbf{x},\mathbf{t})}{\big{[}d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau\big{]}}+\operatorname*{\mathbb{E}}\limits_{\mathbf{x}\sim p(\mathbf{x})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{t}\sim p(\mathbf{t})}\big{[}e^{d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau}\big{]}\Big{]}$ (14)
$+\operatorname*{\mathbb{E}}\limits_{\mathbf{t}\sim p(\mathbf{t})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{x}\sim p(\mathbf{x})}\big{[}e^{d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau}\big{]}\Big{]}.$
@@
This proof is done by extending the proof of Theorem 1 in Wang & Isola (2020), mainly depending on the Continuous Mapping Theorem and the law of large numbers.
$\lim_{N\rightarrow\infty}\mathcal{L}$ $=\lim_{N\rightarrow\infty}\Big{(}-\frac{1}{N}\sum_{i=1}^{N}\log\frac{e^{d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{i})\big{)}/\tau}}{\sum_{j=1}^{N}e^{d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{j})\big{)}/\tau}}-\frac{1}{N}\sum_{i=1}^{N}\log\frac{{e^{d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{i})\big{)}/\tau}}}{\sum_{j=1}^{N}e^{d\big{(}\mathbf{f}_{x}(\mathbf{x}_{j}),\mathbf{f}_{t}(\mathbf{t}_{i})\big{)}/\tau}}\Big{)},$
$=\lim_{N\rightarrow\infty}\Big{(}-\frac{2}{N}\sum_{i=1}^{N}{d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{i})\big{)}/\tau}+\frac{1}{N}\sum_{i=1}^{N}\log{\sum_{j=1}^{N}e^{d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{j})\big{)}/\tau}}+\frac{1}{N}\sum_{i=1}^{N}\log{\sum_{j=1}^{N}e^{d\big{(}\mathbf{f}_{x}(\mathbf{x}_{j}),\mathbf{f}_{t}(\mathbf{t}_{i})\big{)}/\tau}}\Big{)}$
$=\lim_{N\rightarrow\infty}\Big{(}-\frac{2}{N}\sum_{i=1}^{N}{d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{i})\big{)}/\tau}+\frac{1}{N}\sum_{i=1}^{N}\log\frac{1}{N}{\sum_{j=1}^{N}e^{d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{j})\big{)}/\tau}}+\frac{1}{N}\sum_{i=1}^{N}\log\frac{1}{N}{\sum_{j=1}^{N}e^{d\big{(}\mathbf{f}_{x}(\mathbf{x}_{j}),\mathbf{f}_{t}(\mathbf{t}_{i})\big{)}/\tau}}$
$+\frac{2}{N}\sum_{i=1}^{N}\log N\Big{)}$
$=-2\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{x},\mathbf{t})\sim p(\mathbf{x},\mathbf{t})}{\big{[}d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau\big{]}}+\operatorname*{\mathbb{E}}\limits_{\mathbf{x}\sim p(\mathbf{x})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{t}\sim p(\mathbf{t})}\big{[}e^{d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau}\big{]}\Big{]}$
$+\operatorname*{\mathbb{E}}\limits_{\mathbf{t}\sim p(\mathbf{t})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{x}\sim p(\mathbf{x})}\big{[}e^{d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau}\big{]}\Big{]}+2\log N.$
A.3 Relation with Recovering All Information

### v2

@@ 6 Synthetic Experiments and Real-World Evaluation with Pretrained CLIP
Table 1: Assessing identifiability up to linear (a) and permutation (b) transformations under varying assumptions. The first row in (a) and the first two rows in (b) represent settings that align with our assumptions in Corollary 4.2 and Corollary 4.4, while the others show results for violated assumptions. S: Space, Sp: Sphere, U: Uniform, v: vMF ($k=1$), L: Laplace ($\lambda=0.05$), N: Normal ($\delta=0.05$), B: Box, Un: Unbounded, G: GenNorm($\beta=3$).
@@
Table 2: Assessing identifiability up to linear.
@@
Table 3: Assessing identifiability up to permutation.
To evaluate linear identifiability result in Corollary 4.2, we fit a linear regression model between the ground-truth $\mathbf{z}_{x}$ and representations $\mathbf{f}_{x}(\mathbf{x})$ learned by multimodal contrastive learning and report the coefficient of determination ($R^{2}$). Further, to evaluate permutation identifiability result in Corollary 4.4, we employ the mean correlation coefficient (MCC) between the ground-truth $\mathbf{z}_{x}$ and representations $\mathbf{f}_{x}(\mathbf{x})$ learned by multimodal contrastive learning. The first row in Table 3 (a) and the first two rows in Table 3 (b), corresponding to the setting where the assumptions are satisfied, verify the identifiability results on hypersphere and convex bodies, respectively. Our empirical investigations have yielded a critical insight: discrepancies in the assumptions concerning marginal and conditional distributions, as well as the nature of the spaces (hypersphere and convex body), do not significantly impact performance. This robustness is demonstrated by the results detailed in Table 3 (a) for the hypersphere space and Table 3 (b) for convex bodies. This observation is similar to reports from studies conducted in the context of single-model context [67]. This observation might be attributed to the fact that the loss function described in Eq. (2) predominantly relies on the computation of expectations, inherently allowing for a wide range of approximations. If we can approximate the expectation calculations consistently across various distributions and spaces, it might be reasonable that the identifiability results remain well within acceptable bounds.
Real-World Evaluation with Pretrained CLIP
In the context of real data, the true latent coupled variables are unknown. Therefore, we evaluate our theoretical findings from the perspective of disentanglement as discussed in Section 5. Again, we emphasize that, in contrast to previous studies on identifiability for multimodal contrastive learning [8, 62, 14], which rely on simulation experiments, this work validates identifiability through empirical analysis on real-world datasets and pre-trained CLIP model. This underscores the practicality of our theoretical contributions.
@@
According to Section 5, we first extract representations from the pre-trained CLIP model and then apply FastICA to these representations to achieve final representations for CelebA data [38]. We expect these final representations to exhibit clear signs of disentanglement. To validate this, we proceed to train a decoder that reconstructs observational data using these extracted representations. Figure LABEL:fig:_celeba illustrates the effectiveness of our method through latent space traversals. Specifically, it visualizes changes in reconstructions as we traverse one dimension of the latent space at a time, showcasing 4 out of 16 attributes uncovered by our approach. Additional results are available in Appendix H. This achievement not only validates our identifiability results, but also offers a new research line, i.e., learning disentangled representations by CLIP, or exploring how this disentanglement potential relate to the manipulation of pre-trained vision models, such as diffusion models.
Few-shot learning and domain generalization
The goal of disentangled representations is to learn representations that transfer easily and robustly to downstream tasks, making them well-suited for few-shot learning and resilient to distribution shifts [12]. We thus focus on few-shot learning and domain generalization tasks to further evaluate our disentanglement findings.
Figure 6: Quantitative results for 2-shot learning and domain generalization by different methods. ①: Linear Probe, ②: ① with FastICA, and ③: ① with PCA and FastICA.
@@
RN50 31.95 26.48 8.41 20.74 7.44 15.77
34.06 28.74 8.37 21.72 10.15 17.25
34.12 28.68 11.55 25.57 10.15 18.99
RN101 37.64 31.45 13.71 31.09 11.85 20.03
39.58 33.15 13.49 30.29 14.77 22.93
39.86 33.58 17.93 35.48 14.20 25.29
ViT32 38.23 32.00 16.17 33.67 12.88 23.68
40.21 33.97 16.54 34.79 15.72 25.26
39.34 33.44 19.02 36.98 14.69 26.03
ViT16 44.97 38.11 22.06 43.86 25.99 32.51
45.52 39.38 22.55 45.33 30.47 34.43
46.57 40.66 26.67 49.69 31.48 37.13
We extract representations from a limited set of labeled samples using a pre-trained CLIP model, combined with FastICA to align with the hypersphere, and with PCA followed by FastICA to align with convex body. These representations are then used to train a linear classifier. We evaluate the methods on ImageNet [9] for few-shot learning and test robustness on ImageNet-V2 [48], ImageNet-Sketch [59], ImageNet-R [19], and ImageNet-A [20]. Table 6 presents the performance metrics of the proposed methods in few-shot learning (as shown in the ‘SOURCE’ column) and domain generalization (indicated in the ‘TARGET’ columns). Analyzing the data in the SOURCE’ column reveals that the proposed methods outperform the baseline approach, which trains a linear classifier using representations directly obtained from pre-trained CLIP (i.e., the Linear Probe). This superior performance validates our disentanglement findings. Additionally, observations from the TARGET’ column further reinforce the benefits of disentanglement. Refer to Appendix I for more results.
[…]
Leveraging the Disentanglement Potential of CLIP-like Models for Few-Shot Learning
In the final experiments, we demonstrate how the disentanglement potential pushes the boundaries of leveraging pre-trained models, e.g., CLIP. Recent progress shows that pre-trained CLIP’s adaptability can be significantly improved with just a few labeled training samples. The key to leveraging pre-trained CLIP for few-shot learning is effectively utilizing its extracted representations from limited labeled data, as Tip-Adapter and Tip-Adapter-F methods proposed in the work [63]. As we claim, leveraging disentangled potential of pre-trained CLIP can enhance performance on tasks that rely on disentangled representations, including few-shot learning. Therefore, rather than using CLIP’s raw representations, we apply FastICA to extract disentangled representations for few-shot tasks. This can be implemented in a plug-and-play way. As shown in Figure 7, incorporating FastICA in the methods in [63], termed Tip-Adapter with FastICA and Tip-Adapter-F with FastICA, results in better performance, across 11 datasets, including ImageNet [9], Caltech101 [10], FGVCAircraft [40], UCF101 [54], EuroSAT [18], Flowers102 [42], StanfordCars [34], DTD [7], Food101 [3], OxfordPets [44], and, SUN397 [61]. See Appendix J for more details.
@@ 7 Conclusion
In this work, we propose a novel latent partial causal model for multimodal data that moves beyond the traditional DAG structure, using latent coupled variables connected by undirected edges to capture transferable knowledge across modalities. We establish a theoretical link between this generative model and multimodal contrastive learning, showing that the representations learned by multimodal contrastive learning correspond to latent variables in the generative model, with linear and permutation transformations in hypersphere and convex body spaces, respectively. Our results provide the first theoretical guarantees for the disentanglement capabilities of multimodal contrastive learning, with applications in tasks like few-shot learning and domain generalization. Unlike prior simulation-based studies, our work demonstrates the real-world utility of multimodal contrastive learning and offers insights into leveraging pre-trained models like CLIP for disentangled representations. Our model challenges conventional DAG assumptions and provides a flexible, practical framework that enhances the effectiveness of multimodal contrastive learning.
Appendix
[…]
Appendix A Related Work
@@
Multi-modal contrastive representation learning, driven by underlying transferable knowledge across modalities, aims to coalesce inputs from these diverse sources into a cohesive representation space. This is typically achieved using a symmetric version of the standard contrastive loss [43, 16], a method designed to align accurate pairings while distinguishing incorrect ones [17, 47]. Although this approach has proven successful in a range of downstream tasks [47, 65, 66, 39, 2], there remains a gap in our comprehensive theoretical and empirical understanding of the representations it learns. Recently, there has been a growing interest in exploring multi-modal contrastive learning from various perspectives. For instance, the study by [35] provides insights into the modality gap inherent in multi-modal contrastive learning. Similarly, the research presented by [41] establishes a link between general multimodal contrastive loss and SVD analysis. Additionally, [22] posits that learning with multiple modalities can lead to a reduced population risk compared to using a subset of these modalities. Diverging from these approaches, our work delves into multi-modal contrastive representation learning by examining its connection with generative models.
@@
Past research has sought to comprehend the representations derived from standard single-modality contrastive learning, examining them through the lens of alignment and uniformity [60], showing guarantees on the performance of the learned representations on the average classification task [50], or in terms of the identifiability of latent variables [67, 58]. Building on these foundations, our work takes a foreword step. We demonstrate that multi-modal contrastive learning can identify latent coupled variables, extending the insights from previous studies into the realm of multi-modality. Refer to Section G for more details.
@@
Very recently, several studies have emerged, focusing on multi-modal settings [8, 62]. A clear distinction is that: the proposed model captures transferable knowledge across modalities by an undirected edge between latent coupled variables, while previous works often achieve it by introducing shared variables [8, 62]. Notably, our modeling approach is more general, as it can be reduced to the shared variables used in previous works [8, 62] by enforcing an identical mapping on the undirected edge between latent coupled variables. Some of these works have only achieved partial identifiability of coupled variables [8, 62], specifically identifying latent content variables but not latent style variables. In contrast, our work achieves comprehensive identifiability results for all latent coupled variables, offering a deeper level of understanding. Our research also diverges from the approach taken in [14] in two key ways: Firstly, we model ransferable knowledge across modalities using conditional distributions, whereas the latter utilizes identical variables for this purpose. Secondly, while [14] relies on the premise that the mapping from the latent space to observations must be constrained by component-wise corrupters to ensure identifiability, our findings do not necessitate such constraints. Refer to Section F for more details.
@@
Nonlinear Independent Component Analysis (ICA) aims to unravel latent independent variables from observational data that has been subject to a nonlinear mixture of these latent factors. However, as pointed out in the seminal work by [27], solving this problem is generally infeasible without specific underlying assumptions. A prominent direction in contemporary research leverages the concept of distributional changes in latent variables, which leads to the creation of multi-domain observational data. This approach has been extensively explored and developed in a series of studies [25, 26, 29, 32], each contributing to a deeper understanding and more refined methodologies in the field of Nonlinear ICA. We build upon this body of research by incorporating co-occurrence patterns observed across multiple modalities. It is important to note the distinct difference between multi-domain and multi-modal approaches. The former typically implies a consistent mapping from the latent space to the observational space across all domains, whereas the latter accommodates different mappings for each modality. Additionally, while multi-domain approaches generally assume a totally shared latent variables across all domains, multi-modal methods allow for the existence of modality-specific latent variables.
Appendix B The Proof of Theorem 3.1
@@
(The asymptotics of $\mathcal{L}$) For fixed $\tau>0$, as the sample size $N\rightarrow\infty$, the (normalized) multimodal contrastiveloss converges to
$\lim_{N\rightarrow\infty}\mathcal{L}-2\log N=$ $2\operatorname*{\mathbb{E}}\limits_{(\mathbf{x},\mathbf{t})\sim p(\mathbf{x},\mathbf{t})}{\big{[}d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau\big{]}}+\operatorname*{\mathbb{E}}\limits_{\mathbf{x}\sim p(\mathbf{x})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{t}\sim p(\mathbf{t})}\big{[}e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau}\big{]}\Big{]}$ (10)
$+\operatorname*{\mathbb{E}}\limits_{\mathbf{t}\sim p(\mathbf{t})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{x}\sim p(\mathbf{x})}\big{[}e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau}\big{]}\Big{]}.$
@@
This proof is done by mainly depending on the Continuous Mapping Theorem and the law of large numbers.
$\lim_{N\rightarrow\infty}\mathcal{L}-2\log N$ $=\lim_{N\rightarrow\infty}\Big{(}-\frac{1}{N}\sum_{i=1}^{N}\log\frac{e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{i})\big{)}/\tau}}{\sum_{j=1}^{N}e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{j})\big{)}/\tau}}$
$-\frac{1}{N}\sum_{i=1}^{N}\log\frac{{e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{i})\big{)}/\tau}}}{\sum_{j=1}^{N}e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}_{j}),\mathbf{f}_{t}(\mathbf{t}_{i})\big{)}/\tau}}\Big{)}-2\log N,$
$=\lim_{N\rightarrow\infty}\Big{(}\frac{2}{N}\sum_{i=1}^{N}{d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{i})\big{)}/\tau}+\frac{1}{N}\sum_{i=1}^{N}\log{\sum_{j=1}^{N}e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{j})\big{)}/\tau}}$
$+\frac{1}{N}\sum_{i=1}^{N}\log{\sum_{j=1}^{N}e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}_{j}),\mathbf{f}_{t}(\mathbf{t}_{i})\big{)}/\tau}}\Big{)}-2\log N$
$=\lim_{N\rightarrow\infty}\Big{(}\frac{2}{N}\sum_{i=1}^{N}{d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{i})\big{)}/\tau}+\frac{1}{N}\sum_{i=1}^{N}\log\frac{1}{N}{\sum_{j=1}^{N}e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{j})\big{)}/\tau}}$
$+\frac{1}{N}\sum_{i=1}^{N}\log\frac{1}{N}{\sum_{j=1}^{N}e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}_{j}),\mathbf{f}_{t}(\mathbf{t}_{i})\big{)}/\tau}}+\frac{2}{N}\sum_{i=1}^{N}\log N\Big{)}-2\log N$
$=2\operatorname*{\mathbb{E}}\limits_{(\mathbf{x},\mathbf{t})\sim p(\mathbf{x},\mathbf{t})}{\big{[}d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau\big{]}}+\operatorname*{\mathbb{E}}\limits_{\mathbf{x}\sim p(\mathbf{x})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{t}\sim p(\mathbf{t})}\big{[}e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau}\big{]}\Big{]}$
$+\operatorname*{\mathbb{E}}\limits_{\mathbf{t}\sim p(\mathbf{t})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{x}\sim p(\mathbf{x})}\big{[}e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau}\big{]}\Big{]}.$
Appendix C Relation with Recovering All Information